Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 103; σ = 14

P(x ≥ 120) =
Answer #1

Solution :

Given that ,

mean = \mu = 103

standard deviation = \sigma = 14

P(x \geq 120) = 1 - P(x  \leq 120)

= 1 - P[(x - \mu ) / \sigma \leq (120 - 103) / 14]

= 1 -  P(z \leq 1.2143)   

= 1 - 0.8877

= 0.1123

P(x \geq 120) = 0.1123